OpenStudy (anonymous):

i have to find x for this question is it right abc=cbd 3x-30=x+10 +30 +30 3x=x+40 -x -x 2x=40 x=20

OpenStudy (anonymous):

yep good job!

OpenStudy (anonymous):

wicked :D

OpenStudy (anonymous):

To double check you could just plug in the number the you got for x :)

OpenStudy (anonymous):

3(20) -30 = 20 + 10 First multiply 2 x 20 60 - 30 = 30 Then subtract 60 - 30 30 = 30 So you were right! ;)